Davis Division uses two raw materials-X33 and Z42-to produce its product called Base 23, a chemical base for interior house pa. Price b.Quantity C. Mix d. Yield
Answer #1
Standard (100 litres) Standard (18000 litres) Mix Actual (18000 litres)
Qty Rate Amount Qty Rate Amount Quantity Qty Rate Amount
X33 80 $      1.00 $   80.00 14400 $      1.00 $ 14,400.00 13500 16200 $      1.00 $ 16,200.00
Z42 20 $      5.00 $ 100.00 3600 $      5.00 $ 18,000.00 4500 5400 $      5.09 $ 27,500.00
18000 $ 32,400.00 18000 21600 $ 43,700.00
Price Variance = (Standard Price - Actual Price) x Actual Quantity
X33 = ($ 1 - 1) x 16200
Z42 = ($ 5 - 5.09) x 5400 $      500.00
Total $      500.00 (Unfavorable)
Quantity Variance = (Standard Quantity - Actual Quantity) x Standard Price
X33 = (14400 - 16200) x $ 1 $   1,800.00 (Unfavorable)
Z42 = (3600 - 5400) x $ 5 $   9,000.00 (Unfavorable)
Total $ 10,800.00 (Unfavorable)
Mix Variance = (Mix Quantity - Actual Quantity) x Standard Price
X33 = (13500 - 16200) x $ 1 $   2,700.00 (Unfavorable)
Z42 = (4500 - 5400) x $ 5 $   4,500.00 (Unfavorable)
Total $   7,200.00 (Unfavorable)
Yield Variance = (Standard Quantity - Mix Quantity) x Standard Price
X33 = (14400 - 13500) x $ 1 $      900.00 (favorable)
Z42 (3600 - 4500) x $ 5 $   4,500.00 (Unfavorable)
Total $   3,600.00 (Unfavorable)